ABC Pvt Ltd proposes to buy a truck that costs š¯‘…š¯‘€ 50,000. The company has two alternatives:


Alternative I: Buy from Nice Car Ltd by making a down payment of š¯‘…š¯‘€ 15,000 and settling the balance with 60 monthly payments at 10 % per annum flat rate.


Alternative II: Buy from Cheap Car Ltd by making a down payment of š¯‘…š¯‘€ 10,000 and settling the balance with 60 monthly payments of š¯‘…š¯‘€ 690 each.


Which alternative should the company select? Show it by calculation in detail for both alternatives.


1
Expert's answer
2021-02-25T01:51:46-0500

The company should choose alternative 2 because it has low interest of Rm 1200.

The first alternative interest

"=5000-15000=35000"

"=60*0.01\/12*35000=17500"

"=35000+17500=52500"

Interest will be

"=52500-5000"

"=Rm 2500"

The second alternative interest

"=60*690"

"=Rm41400"

Interest will be

"=50000-10000=40000"

"=41400-40000= Rm 1400"